Assignment Number 5   MELBOURNE ENGINEERING

Shiela Forbes is the Chief Financial Officer [CFO] of MELBOURNE ENGINEERING, a U.S. based manufacturer of gas turbine equipment. She has just concluded negotiations for the sale of a turbine generator to Crown, a British firm for THREE million pounds. This single sale is quite large in relation to MELBOURNE ENGINEERING’s present business. MELBOURNE ENGINEERING has no other current foreign customers, so the currency risk of this sale is of particular concern. The sale is made in December with payment due six months later in June. Shiela Forbes has collected the following financial market information for the analysis of her currency exposure problem:

  • Spot Exchange rate: $1.2940 per British pound.
  • Six-month forward rate: $1.2949 per pound
  • MELBOURNE MANUFACTURING’s cost of capital: 14%
  • U.K. annual borrowing interest rate: 16.0% (or 8% interest for six months)
  • U.K. annual investment interest rate: 14.0% (7% interest for six months)
  • U.S. annual borrowing interest rate: 8.0%     (4.0% interest for six months)
  • U.S. one-year investment interest rate: 6.0% (3% interest for six months)
  • June put option in the over-the-counter (bank) market for 1,000,000 British pounds;

Strike price $1.27 (nearly at-the money) 2.0% premium

  • MELBOURNE MANUFACTURING’s advisory service forecast the spot rate in six months will be $1.28/pound.

Like many manufacturing firms, MELBOURNE MANUFACTURING operates on relatively narrow margins. Although Ms. Forbes and MELBOURNE MANUFACTURING would be very happy if the pound appreciated versus the dollars, concerns center on the possibility that the pound will fall. When Ms. Forbes budgeted this specific contract, she determined that the minimum acceptable margin was at a sale price of $3,750,000. The budget rate, the lowest acceptable dollar per pound exchange rate, was therefore established at $1.25 per British pound. Any exchange rate below would result in MELBOURNE ENGINEERING actually losing money on the transaction.

Four alternatives are available to MELBOURNE ENGINEERING to manage the exposure:

  1. Remain un-hedged.
  2. Hedge in the forward market.
  3. Hedge in the money market.
  4. Hedge in the options market.    What should MELBOURNE ENGINEERING do?
Answer #1

A. Remain un-hedged:

Expected amount to be received at the end of six months:

Expected spot rate after six months =$1.28/Pound

Expected receipt in dollars after six months=1.28*3 million=$3,840,000

B.Hedge in the forward market

Six month Forward rate =$1.2949

Expected receipt in dollars after six months=1.2949*3 million=$3,884,700

C Hedge in the money market:

Borrow (3million/1.08) Pounds in UK at 16% interest

Amount borrowed  in Pounds in UK =3million/1.08=2,777,778 Pounds

Future value of this amount after six months will be 3 million pounds which can be paid from the receivable.

Convert 2,777,778 pounds into dollar at current spot rate $1.2940 per pound

Amount of dollar available after conversion=1.2940*2777778=$3,594,444

Invest $3,594,444 in US at 6%

Future Value after six months =$3,594,444*1.03=$3,702,278

D. Hedge in option market:

Strike Price =$1.27 per pound

Cost of Option =2%*1.27*3 million =$76,200

If the exchange rate falls below $1.27, there will be gain in option which will offset loss in spot market

Net amount to be received =Amount as per spot rate -Cost of option

Net amount to be received =1.28*3 million -76200=$3,840,000-$76,200=$3,763,800

EXPECTED AMOUNT TO BE RECEIVED AT END OF SIX MONTHS:

A. Remain un-hedged:$3,840,000

B.Hedge in the forward market:$3,884,700

C Hedge in the money market:$3,702,278

D. Hedge in option market:$3,763,800

What should MELBOURNE ENGINEERING do?

Since the forecast spot rate is $1.28/Pound, Melbourne Engg should Hedge in the forward market:

ANSWER:

B.Hedge in the forward market: